Ergonomics Inc. sells ergonomically designed office chairs. The company has the following information:

Average demand = 22 units per day

Average lead time = 31 days

Item unit cost = $51 for orders of less than 210 units

Item unit cost = $47 for orders of 210 units or more

Ordering cost = $26

Inventory carrying cost = 20%

The business year is 250 days

Assume there is no uncertainty at all about the demand or the lead time.

a. Calculate EOQ if unit cost is $51 and $47. (Note: These EOQs do not need to be feasible in their price range.) (Round up your answers to the next whole number.)

b. Calculate annual ordering costs for each alternative? (Round your answers to 2 decimal places.)

c. Calculate annual inventory carrying costs for each alternative? (Round your answers to 2 decimal places.)

d. Calculate annual product costs for each alternative?

e. What will be the total costs for each alternative? (Round your answers to 2 decimal places.)

f. Based on your analysis, how many chairs should they order at a time? (Round your answers to 2 decimal places.)

   

g. How much the firm can save annually by using the order quantity in Part f. instead of the first EOQ shown in Part a? (Round your answer to 2 decimal places.)

*Please solve all questions*
Answer #1

Average demand, d = 22 units per day

Average lead time, L = 31 days

Item unit cost, C1 = $51 for orders of less than 210 units

Item unit cost, C2 = $47 for orders of 210 units or more

Ordering cost, S = $26

Inventory carrying cost, h = 20%

The business year is 250 days

Annual demand, D = 250*22 = 5500

a)

For unit cost of $ 51, carrying cost, H = 51*20% = 10.2,

EOQ = SQRT(2DS/H) = SQRT(2*5500*26/10.2) = 168

For unit cost of $ 47, carrying cost, H = 47*20% = 9.4,

EOQ = SQRT(2DS/H) = SQRT(2*5500*26/9.4) = 175 (this is not feasible in its price range, because it is less than 210)

b)

For unit cost of $ 51, Annual ordering cost = (5500/168)*26 = $ 851.19

For unit cost of $ 47, Annual ordering cost = (5500/210)*26 = $ 680.95

c)

For unit cost of $ 51, Annual carrying cost = (168/2)*10.2 = $ 856.80

For unit cost of $ 47, Annual carrying cost = (210/2)*9.4 = $ 987.00

d)

For unit cost of $ 51, Annual product cost = 5500*51 = $ 280,500

For unit cost of $ 47, Annual product cost = 5500*47 = $ 258,500

e)

For unit cost of $ 51, total annual cost = 851.19+856.80+280500 = $ 282,208.00

For unit cost of $ 47, total annual cost = 680.95+987+258500 = $ 260,167.95

f) Total annual cost of $ 47 unit cost is lower. Therefore, 210 chairs should be ordered at a time.

g) Cost saving = 282208 - 260167.95 = $ 22,040.04